OpenStudy (anonymous):

The density of 3 M solution of Nacl is 1.25g / L calculate the molality of the solution?

OpenStudy (ujjwal):

gm/liter=molarity * molecular weight

OpenStudy (anonymous):

We r given all the three

OpenStudy (anonymous):

@Vincent-Lyon.Fr Plz help me!

OpenStudy (ujjwal):

oh, it is asking for molality.

mathslover (mathslover):

Hmn .. Class XI some basic concepts of chemistry

mathslover (mathslover):

3M soln=175.5g in 1000ml of soln right ?

OpenStudy (anonymous):

yup

mathslover (mathslover):

i.e 175.5 g in 1250 g of soln or 3 moles in1074.5 of solvent or 2.79 no of moles in 1000g of solvent. right ?

OpenStudy (anonymous):

ok......... if the density is 1.25 g /ml then

OpenStudy (anonymous):

also the answer is same!!

mathslover (mathslover):

Sorry i am having exams cant help u now BUT i am sure that that will help u

OpenStudy (anonymous):

@mathslover u see my question is density is 1.25 g /L

OpenStudy (anonymous):

not ml....

mathslover (mathslover):

Why not to just put 1.75 instead of 1.25 ?

OpenStudy (anonymous):

hey i am serious!

OpenStudy (anonymous):

PLzz

OpenStudy (anonymous):

The density of 3 M solution of Nacl is 1.25g / L calculate the molality of the solution?

OpenStudy (ujjwal):

Do you have the answer?

OpenStudy (anonymous):

nop!

mathslover (mathslover):

@ujjwal can u help him further with the basis of the image ( soln ) of the simiar question given by me ?

OpenStudy (anonymous):

if the density is 1.25g/ ml then the answer is 2.79

OpenStudy (anonymous):

but my ? asks for 1.25 g /l

OpenStudy (ujjwal):

How do you do it when the density is 1.25g/ml?

OpenStudy (ajprincess):

convert l into ml

OpenStudy (anonymous):

by using maths lover attached a photo

OpenStudy (ujjwal):

So, put the density 1.25 \(\times10^{-3}\)g/ml and then solve it!

OpenStudy (anonymous):

thxxx

OpenStudy (ujjwal):

Well, I have this formula too. May be it will help you too.. \[\textrm{molality}=\frac{\textrm{wt. of solute}\times 1000}{\textrm{mol. wt.}\times\textrm{wt. of solvent (gm)}}\]

OpenStudy (vincent-lyon.fr):

@mathslover gave the right answer, considering the typo in the question. The density of an aqueous solution must be around 1 (non dimensional) That means the volumetric mass is around 1000 kg/m³ = 1000 g/L = 1000 mg/mL, whichever you prefer. 1.25 g/L roughly corresponds to air in standard conditions !!